# Revisiting Quantization of Gauge Field Theories: Sandwich Quantization Scheme

### v2

@@ 3.2 Solving sandwich conditions
To solve (8), we employ the ideas developed in [19, 20]. Consider a Hermitian operator $\bm{\mathcal{C}}$ that acts on a total Hilbert space $\mathcal{H}_{\text{t}}$ with the requirement that all states in $\mathcal{H}_{\text{p}}\subset\mathcal{H}_{\text{t}}$ should satisfy $\langle\psi^{\prime}|\bm{\mathcal{C}}|\psi\rangle=0$. Let $|C\rangle$ denote eigenstates of $\bm{\mathcal{C}}$,
@@
Zero-eigenstates with ${C}=0$ are physical states; however, the sandwich condition $\langle\psi^{\prime}|\bm{\mathcal{C}}|\psi\rangle=0$ allows for other solutions with ${C}\neq 0$.
Assume that $\bm{\mathcal{C}}$ has a $\mathbb{Z}_{2}$ symmetric spectrum such that for every state with eigenvalue $|{C}|$ there is another state with eigenvalue $-|{C}|$. That is, assume there exists a (anti)Hermitian operator þ such that
þ$\textbf{\th}\bm{\mathcal{C}}\textbf{\th}=-\bm{\mathcal{C}},\qquad\textbf{\th}^{2}=1,\qquad\textbf{\th}|C\rangle=|-C\rangle\,.$ (16)
@@
$\begin{split}&\textbf{(A):}\qquad\bm{\mathcal{C}}|\psi\rangle\in\mathcal{H}_{\text{p}},\ \forall|\psi\rangle\in\mathcal{H}_{\text{p}},\\ &\textbf{(B):}\qquad\bm{\mathcal{C}}|\psi\rangle\in\mathcal{H}_{\text{c}},\ \forall|\psi\rangle\in\mathcal{H}_{\text{p}},\quad\mathcal{H}_{\text{c}}:=\mathcal{H}_{\text{t}}-\mathcal{H}_{\text{p}}.\end{split}$ (19)
That is, for (A) case, $\bm{\mathcal{C}}$ is an operator defined over the physical Hilbert space $\mathcal{H}_{\text{p}}$ and hence (8) implies that physical states must be zero eigenstates of $\bm{\mathcal{C}}$. For (B), however, $\bm{\mathcal{C}}$ is defined over the total Hilbert space $\mathcal{H}_{\text{t}}$ and not $\mathcal{H}_{\text{p}}$. In our notation, total Hilbert space is divided into two parts, $\mathcal{H}_{\text{p}}$ and its complement $\mathcal{H}_{\text{c}}$: $\mathcal{H}_{\text{t}}=\mathcal{H}_{\text{p}}\cup\mathcal{H}_{\text{c}}$. Without loss of generality one can choose
@@
Note that (20), within our construction, is an immediate outcome and we need not an independent proof: If $\mathcal{H}_{\text{p}}$ is spanned by $|\mathtt{C}\rangle_{+}$, then $\mathcal{H}_{\text{c}}$ is spanned by $|\mathtt{C}\rangle_{-}$, which by construction are orthogonal sets. [^7] We thank Marco Serone for a question in this regard.
Equipped with the above, we can readily solve (8):
@@
That is, Class 1 Hilbert space is the part of $\mathcal{H}_{\text{t}}$ that is invariant under þ, whereas Class 2 states under þ are mapped onto states in $\mathcal{H}_{\text{c}}$. In other words, $\bm{\mathcal{C}}$ maps a Class 2 physical state onto a state in the complement Hilbert space $\mathcal{H}_{\text{c}}$ while $\bm{\mathcal{C}}^{2}$ takes a physical Class 2 state to a physical state. As pointed out in [19], Class 1 and Class 2 are two super-selection sectors in the physical Hilbert space defined through the quantization scheme in (8). There is no overlap between the Class 1 and Class 2 Hilbert spaces and the Class 1 Hilbert space is a part of $\mathcal{H}_{\text{c}}$ as seen by the Class 2 physical Hilbert space.
We close this section with some important remarks: The vacuum state of the total Hilbert space $\mathcal{H}_{\text{t}}$ is, by definition, a state in Class 1 Hilbert space, and hence cannot be a state in Class 2 Hilbert space. As we will demonstrate in an illustrative example in the next section, Class 2 physical Hilbert space can be constructed by the action of gauge-invariant operators on a Class 2 vacuum state. Our construction above specifies this Class 2 vacuum state. As our example clarifies, there are, in fact, infinitely many such Class 2 vacuum states. The difference between these different physical Hilbert spaces is that they are built upon different vacuum states. That is, all states in Class 1 and 2 Hilbert spaces are made from gauge-invariant operators and we have usual operator-state correspondence (see footnote 1) for all of these physical Hilbert spaces. Identity operator belongs to Class 1 $\mathcal{H}_{\text{p}}$ and is not a physical operator in Class 2. Therefore, if we require that the identity operator should be a part of the physical operators, e.g. as we do in the usual computation of the S-matrix, then we should choose Class 1 Hilbert space.
@@ 4 Example: Maxwell theory
As a very simple but illustrative example, we work through constructing Class 2 states for the $d$ dimensional Maxwell theory, where $\Phi$ are the gauge field components $A_{\mu}$ with the Lagrangian
@@
As we see $\Pi^{t}$ identically vanishes and hence $A_{t}$ is the gauge d.o.f. So, in our notation, $\varphi_{i}$ is $A_{t}$ and $\psi_{A}$ are the spatial components of the gauge field $A_{a}$ and under generic gauge transformations, $A_{\mu}\to A_{\mu}+\partial_{\mu}\lambda$. Note that both $A_{t}$ and $A_{a}$ transform under a gauge transformation. There are various gauges used in the literature, e.g. Lorenz gauge $\partial^{\mu}A_{\mu}=0$ or temporal gauge $A_{t}=\varphi_{0}(\vec{x})$ for a given function $\varphi_{0}(\vec{x})$. Here we adopt the latter, yielding
@@
With the above we have
@@
Note that the consistency of Maxwell’s equations implies $\partial_{t}(\vec{\nabla}\cdot\vec{E})=0$ or $\vec{\nabla}\cdot\vec{E}$ should be constant in time.
The constraint (EoM for $A_{t}$) is the Gauss law: $\vec{\nabla}\cdot\vec{E}=0$, $E^{a}=F^{ta}$ is the electric field strength, the momentum conjugate to $A_{a}$. In this case and since we are dealing with an Abelian gauge theory, the constraint is gauge-invariant (instead of being covariant). Solutions to the classical equations of motion after fixing $A_{t}=\varphi_{0}(\vec{x})$ gauge and imposing the constraint $\vec{\nabla}\cdot\vec{E}=0$ is
@@
The total Hilbert space is now obtained by the action of $(\bm{\epsilon}^{\perp}_{a}(k))^{\dagger},\bm{\varphi_{0}}^{\dagger}(\vec{k})$ on the vacuum state. The constraint is now $\vec{\nabla}\cdot\vec{\mathbf{E}}=0$. Its spectrum exhibits the $\mathbb{Z}_{2}$ symmetry we required in our construction: the operator þ is the charge conjugation. So, we can simply work through the general procedure outlined in the previous section. One readily sees that Class 1 states are the usual transverse photon states we find in QED.
@@
is the observer’s (background) electric charge density operator. For explicit construction of Class 2 states, we use eigenstates of the Hermitian operator $\bm{\mathcal{Q_{B}}}$. Consider eigenstates of $\bm{\mathcal{Q_{B}}}$:
@@
solve the constraint equation. Here $\mathcal{Q_{B}}=\mathcal{Q_{B}}(\vec{x})$ is generic function up to the constraint that $\mathcal{Q_{B}}$ and $-\mathcal{Q_{B}}$ are not viewed as independent. $\mathcal{Q_{B}}=0$ sector is special, as in this case þ$|0\rangle_{+}=0$; it produces Class 1 states. $\mathcal{Q_{B}}\neq 0$ produce Class 2 states.
@@
    Note that we are considering Maxwell’s theory in the absence of electromagnetic genuine physical current and $\bm{\mathcal{Q_{B}}}$ should not be viewed as a genuine (physical) electric charge density operator. We are requiring vanishing of $\bm{\mathcal{Q_{B}}}$ through sandwich conditions.
@@
    We crucially note again that $\bm{\mathcal{Q_{B}}}$ on any Class 2 physical state is not a physical state, it is in the complement part $\mathcal{H}_{\text{c}}$.
@@
  • (6)
    For generic $\mathcal{Q_{B}}(\vec{x})$ in (35), the vacuum state breaks translation symmetry. If we are interested in translation-symmetric vacua, one may restrict the Class 2 states to a constant, non-negative $\mathcal{Q_{B}}$ sector.
@@
[…]
@@
    $|\mathcal{Q_{B}}\rangle_{+}$ is the sum of two eigenstates of $\mathcal{Q_{B}}$ with opposite charge densities. Thus, in a loose sense, one may think of these states as having a vanishing “total electric charge”. For the special super-selection sector with constant ($x$-independent) $\mathcal{Q_{B}}$, there is an “electric dipole” associated with the vacuum state.
  • (10)
    Charge conjugation is not an operator well-defined over a Class 2 physical Hilbert space $\mathcal{H}_{\text{p}}$. However, it is physically expected that charge conjugation invariance should be retained for physical observables within QED, in which physical observables are of the form of gauge-invariant operators.
@@
    Since the theory is gauge invariant, it is expected that these super-selection sectors in Class 2 Hilbert space do not mix with each other. This point should be more rigorously verified. We postpone this to future studies.
@@ 5 Outlook
In this short note, we revisited the old problem of the quantization of gauge field theories. We argued that there is a small, but potentially important point that has slipped the attention of physicists. We stressed the already known fact that (e.g. see [14, 17, 18]) for quantization it is sufficient to impose the EoM for the gauge d.o.f, which appear as constraints, through sandwich conditions (8). We briefly discussed how the sandwich conditions may be solved, leading to a new set of super-selection sectors in physical Hilbert space besides the standard Class 1 states.
Class 1 physical Hilbert space, obtained by imposing gauge-fixing constraints through the right-action (6), is a direct extension of the classical constraints to the quantum level. The sandwich conditions and Class 2 solutions are the options arising through the quantization procedure, with no direct classical counterpart. The latter may be witnessed by the appearance of $\hbar$ in the right-hand side of (13). At the quantum level, we have the option that while $\bm{\mathcal{C}}^{i}$ vanish when inserted into the $n$-point function of any generic local operators, the operator $\bm{\mathcal{C}}^{i}$ need not be well-defined over the physical Hilbert space $\mathcal{H}_{\text{p}}$. In other words, requiring that $\bm{\mathcal{C}}^{i}$ are well defined over $\mathcal{H}_{\text{p}}$ restricts us to Class 1 Hilbert space, whereas relaxing this seemingly unnecessary requirement, we have the option of super-selection sectors in Class 2 Hilbert space. Some steps are remaining to complete the sandwich quantization proposal and show its consistency and sufficiency.
@@
In this work, for illustration purposes we focused on the “axial-type gauges”. Consistency of the proposal requires that it should work for any arbitrary gauge-fixing. For these cases one should work through introducing ghosts and check the BRST invariance, that the physical states satisfying the sandwich conditions are defined up to BRST equivalence classes (see e.g. [6] for discussion on BRST symmetry). That is, we need to show that our Class 2 Hilbert space, and in fact each super-selection sector thereof, is BRST invariant.
@@
We need to uncover the physical meaning of the new sectors in the physical Hilbert space, the super-selection sectors in the Class 2 part. First we note that, Class 2 states by definition, are orthogonal to Class 1 states, and the gauge-invariant dynamics of the theory will never mix them with Class 1 states. That is, one can consistently formulate the theory using only Class 1 states and recover the standard QFT textbook results. There is, however, another option: physics may be formulated based on either of the super-selection sectors in Class 2 Hilbert space. Based on the example explored in [20] and the Maxwell theory example discussed in the previous section, we propose the “sandwich equivalence principle”:
Physical observables of gauge field theories should equivalently be described by Class 1 or either of super-selection sectors in Class 2 physical Hilbert spaces.
The above is somewhat similar to what we have in Einstein’s Equivalence Principle and the choice of different observers: each super-selection sector in Class 2 physical Hilbert space provides states above a vacuum state associated with a particular physical observer. Of course, it remains to fully establish this equivalence principle.
We close by mentioning that sandwich conditions were discussed in (very) early QED literature, see e.g. [13, 14] or in the string textbook, see first page of chapter 4 in [18]. To the knowledge of the author, the reincarnation of sandwich quantization scheme in more recent literature happened with a different motivation and argument than the one presented here. It was first presented for the quantization of null strings, i.e. worldsheet theory of tensionless strings whose worldsheet is a 2d null surface [21]. It was then noted that the same quantization scheme may be invoked for generic string worldsheet theory [20] and also null $p$-brane theory [19]. In this note we put these previous examples in a broader framework: sandwich quantization scheme can/should be invoked for any field theory with local (gauge) symmetry.

### v3

@@ 3.2 Solving sandwich conditions
To solve (8), we first employ the ideas developed in [8, 2] and then use the construction for the case of a field theory.
General idea of the solution.
Consider a Hermitian operator $\bm{\mathcal{C}}$ that acts on the total vector space $\mathcal{H}_{\text{tot}}$ with the requirement that all states in $\mathcal{H}_{\text{p}}\subset\mathcal{H}_{\text{tot}}$ should satisfy $\langle\psi^{\prime}|\bm{\mathcal{C}}|\psi\rangle=0$. Let $|C\rangle$ denote eigenstates of $\bm{\mathcal{C}}$,
@@
Zero-eigenstates with C=0{C}=0 are physical states; however, the sandwich condition $\langle\psi^{\prime}|\bm{\mathcal{C}}|\psi\rangle=0$ allows for other solutions with ${C}\neq 0$.
Assume that $\bm{\mathcal{C}}$ has a $\mathbb{Z}_{2}$ symmetric spectrum such that for every state with eigenvalue |C||{C}| there is another state with eigenvalue |C|-|{C}|. That is, assume there exists a (anti)Hermitian operator þ such that
$\textbf{\th }\bm{\mathcal{C}}\textbf{\th }=-\bm{\mathcal{C}},\qquad\textbf{\th }^{2}=1,\qquad\textbf{\th }|C\rangle=|-C\rangle\,.$ (16)
@@
$\begin{split}&\textbf{(A):}\qquad\bm{\mathcal{C}}|\psi\rangle\in\mathcal{H}_{\text{p}},\ \forall|\psi\rangle\in\mathcal{H}_{\text{p}},\\ &\textbf{(B):}\qquad\bm{\mathcal{C}}|\psi\rangle\in\mathcal{H}_{\text{c}},\ \forall|\psi\rangle\in\mathcal{H}_{\text{p}},\quad\mathcal{H}_{\text{c}}:=\mathcal{H}_{\text{tot}}-\mathcal{H}_{\text{p}}.\end{split}$ (19)
That is, for the case (A), $\bm{\mathcal{C}}$ is an operator defined over the physical Hilbert space $\mathcal{H}_{\text{p}}$ and hence (8) implies that physical states must be zero eigenstates of $\bm{\mathcal{C}}$. For (B), however, $\bm{\mathcal{C}}$ is defined over $\mathcal{H}_{\text{tot}}$ and not $\mathcal{H}_{\text{p}}$. In our notation, $\mathcal{H}_{\text{tot}}$ is divided into two parts, $\mathcal{H}_{\text{p}}$ and its complement $\mathcal{H}_{\text{c}}$: $\mathcal{H}_{\text{tot}}=\mathcal{H}_{\text{p}}\cup\mathcal{H}_{\text{c}}$. Without loss of generality one can choose
@@
Note that within our construction, (20) is an immediate outcome and we need not an independent proof: If $\mathcal{H}_{\text{p}}$ is spanned by $|\mathtt{C}\rangle_{+}$, then $\mathcal{H}_{\text{c}}$ is spanned by $|\mathtt{C}\rangle_{-}$, which are by construction orthogonal sets. [^8] We thank Marco Serone for a question in this regard.
Application to the case of a gauge field theory.
The above was a general construction which works well (without extra mathematical subtleties) for finite or non-continuous infinite-dimensional Hilbert spaces. For the case of quantum field theories, where states in $\mathcal{H}_{\text{tot}}$ have a continuous label, one should be more careful with some mathematical subtleties. Here, we do not intend to delve into and address all those mathematical subtleties. They need a more in-depth study and will distract the reader from the main conceptual points we intend to raise and discuss in this short and educational note. They will be reported upon in future publications. The first issue is with (15). One should note that for a generic operator-function $\bm{\mathcal{C}}$, (15) is only meaningful if $\bm{\mathcal{C}}$ at different points in space commute with each other. [^9] We thank Sergio Cecotti for discussion on this issue. In field theories where we usually make a perturbative analysis, one can work with Fourier transform of the constraint $\bm{\mathcal{C}}$, and this issue requires commuting constraints at different values of momenta. Moreover, as in the usual Schwinger-Gupta-Bleuler method, we only need to consider (15) for positive frequency modes or in cases like that of string theory [2] where we deal with constraints quadratic in creation-annihilation operators, we consider the normal ordered constraints. More on discussions on these subtleties will be given below. For now and with these precautions and just to illustrate the basic idea, let us employ the above $\mathbb{Z}_{2}$ based construction, to solve (8):
@@
That is, Class 1 Hilbert space is the part of $\mathcal{H}_{\text{tot}}$ that is invariant under þ, whereas Class 2 states under þ are mapped onto states in $\mathcal{H}_{\text{c}}$. In other words, $\bm{\mathcal{C}}$ maps a Class 2 physical state onto a state in the complement Hilbert space $\mathcal{H}_{\text{c}}$ while $\bm{\mathcal{C}}^{2}$ takes a physical Class 2 state to a physical state. As pointed out in [8], Class 1 and Class 2 are two super-selection sectors in the physical Hilbert space defined through the quantization scheme in (8). There is no overlap between the Class 1 and Class 2 Hilbert spaces and the Class 1 Hilbert space is a part of $\mathcal{H}_{\text{c}}$ as seen by the Class 2 physical Hilbert space.
We close this section with some important remarks:
  1. 1.
    The vacuum state of the total Hilbert space $\mathcal{H}_{\text{tot}}$ is, by definition, a state in the Class 1 Hilbert space, and hence cannot be a state in the Class 2 Hilbert space.
  2. 2.
    The difference between Class 1 and Class 2 Hilbert spaces is in the vacuum state upon which they are built. Explicitly, as is textbook material, all states in the Class 1 physical Hilbert space can be constructed by the action of all gauge-invariant local operators on the Class 1 vacuum state. In the OSC terminology, the Class 1 vacuum state corresponds to the identity operator in this physical Hilbert space.
  3. 3.
    In a similar way, the Class 2 physical Hilbert space is constructed by the action of the same gauge-invariant local operators on the Class 2 vacuum state.
  4. 4.
    However, the crucial fact to note is that Class 2 Hilbert space consists of infinitely many (a continuum for a generic gauge field theory) of super-selection sectors. That is, we are not dealing with a single Class 2 vacuum state or Class 2 physical Hilbert space; we have a continuum of Class 2 vacuum states or physical Hilbert spaces.
  5. 5.
    To understand the above comment, recall the discussion above (21) and that the constraint is an operator on the spacetime. One can require the Class 2 vacuum state corresponds to a non-vanishing Fourier mode of the constraint for a given momentum. We will make this a bit more explicit in the example we discuss in the next section.
  6. 6.
    In a different wording, the difference between different physical Hilbert spaces, Class 1 or Class 2 cases, is that they are built upon different vacuum states. Each super-selection sector has its own “identity operator,” nonetheless, these Hilbert spaces are in one-to-one correspondence and the consistency of construction requires that physics built upon either of these Hilbert spaces should be the same. So, working with the textbook choice of Class 1 vacuum and Hilbert space is enough to describe physics.
@@ 4 Example: Maxwell theory
As a simple but illustrative example, we work through constructing the Class 2 states for the $d$ dimensional Maxwell theory, where $\Phi$ are the gauge field components $A_{\mu}$ with the Lagrangian
@@
As we see $\Pi^{t}$ identically vanishes and hence $A_{t}$ is the gauge d.o.f. So, in our notation, $\varphi_{i}$ is $A_{t}$ and $\psi_{A}$ are the spatial components of the gauge field $A_{a}$ and under generic gauge transformations, $A_{\mu}\to A_{\mu}+\partial_{\mu}\lambda$. Note that both $A_{t}$ and $A_{a}$ transform under a gauge transformation. There are various gauges used in the literature [14, 10], e.g. Lorenz gauge $\partial^{\mu}A_{\mu}=0$ or temporal gauge $A_{t}=\varphi_{0}(\vec{x})$ for a given function $\varphi_{0}(\vec{x})$. Here we adopt the latter, yielding
@@
With the above, we have
@@
Note that the consistency of Maxwell’s equations implies
$\partial_{t}(\vec{\nabla}\cdot\vec{E})=0\qquad\text{or}\qquad\vec{\nabla}\cdot\vec{E}\ \text{should be constant in time}.$ (31)
The constraint (EoM for $A_{t}$) is the Gauss law: $\vec{\nabla}\cdot\vec{E}=0$, $E^{a}=F^{ta}$ is the electric field strength, the momentum conjugate to $A_{a}$. In this case and since we are dealing with an Abelian gauge theory, the constraint is gauge-invariant (instead of being covariant).
Solutions to the classical equations of motion after fixing $A_{t}=\varphi_{0}(\vec{x})$ gauge and imposing the constraint $\vec{\nabla}\cdot\vec{E}=0$ is
@@
The total vector space is obtained by the action of $(\bm{\epsilon}^{\perp}_{a}(k))^{\dagger},\bm{\varphi_{0}}^{\dagger}(\vec{k})$ on the vacuum state. The constraint is now $\vec{\nabla}\cdot\vec{\mathbf{E}}=0$. Its spectrum exhibits the $\mathbb{Z}_{2}$ symmetry we required in our construction: the operator þ is the charge conjugation. So, we can simply work through the general procedure outlined in the previous section. One readily sees that Class 1 states are the usual transverse photon states we find in QED.
@@
[…]
@@
solve the constraint equation. Here $\mathcal{Q_{B}}=\mathcal{Q_{B}}(\vec{x})$ is generic function up to the constraint that $\mathcal{Q_{B}}$ and $\mathcal{Q_{B}}$ are not viewed as independent. $\mathcal{Q_{B}}=0$ sector is special, as in this case $|0\rangle_{+}=0$; it produces Class 1 states. $\mathcal{Q_{B}}\neq 0$ produce Class 2 states.
@@
[…]
@@
[…]
@@
    For generic $\mathcal{Q_{B}}(\vec{x})$ in (36), the vacuum state breaks translation symmetry. If we are interested in translation-symmetric vacua, one may restrict either to Class 1 or to the Class 2 states with a constant, non-negative $\mathcal{Q_{B}}$ sector. That is, as if the vacuum state has a given uniform constant label (which is not a charge) in the background; see the comment below.
  • (6)
@@
    $|\mathcal{Q_{B}}\rangle_{+}$ is the sum of two eigenstates of $\mathcal{Q_{B}}$ with opposite charge densities. Thus, in a loose sense, one may think of these states as having a vanishing “total electric charge.” For the special super-selection sector with constant ($x$-independent) $\mathcal{Q_{B}}$, there is an “electric dipole” associated with the vacuum state.
@@
    Charge conjugation is not an operator well-defined over a Class 2 physical Hilbert space $\mathcal{H}_{\text{p}}$. However, it is physically expected that charge conjugation invariance should be retained for physical observables within QED, in which physical observables are of the form of gauge-invariant local operators. An explicit demonstration of this fact follows directly from arguments and analysis similarly to those in section 3.1 that led to (14).
@@
    Since the theory is gauge invariant, it is expected that these super-selection sectors in a Class 2 Hilbert space do not mix with each other. This point should be more rigorously verified. We postpone this to future studies.
  • (12)
[…]
@@ 5 Outlook
In this short note, we revisited the old problem of the quantization of gauge field theories. We argued that there is a small, but potentially important point that has slipped the attention of physicists. The sandwich conditions were discussed in (very) early QED literature, see e.g. [9, 5] or in the string textbook, see first page of chapter 4 in [15]. However, it seems to be overlooked or forgotten in the development of quantum gauge field theories. To the knowledge of the author, the reincarnation of sandwich quantization scheme in more recent literature happened with a different motivation and argument than the one presented here. It was first presented for the quantization of null strings, i.e. worldsheet theory of tensionless strings whose worldsheet is a 2d null surface [1]. It was then noted that the same quantization scheme may be invoked for generic string worldsheet theory [2] and also null $p$-brane theory [8]. In this note we put these previous examples in a broader framework: sandwich quantization scheme can/should be invoked for any field theory with local (gauge) symmetry.
Our main motivation for revisiting the quantization of gauge theories, while in part has been better understanding of null strings or branes, has to do with the more fundamental and abstruse problem of the role of observers in quantum theory. This problem has been there since the introduction of quantum mechanics a century ago and has continued in quantum field theory and up to date. There has been a resurgence of the same conceptual issue in the context of quantum gravity and (quantum) cosmology. In dealing with usual gauge theories (like QED, QCD or standard model of particle physics) and as far as computation of $n$-point function of local gauge-invariant operators are concerned, sandwich quantization, as we discussed, seems to be more like an academic question does not add to the computed physical observables. However, in my opinion and understanding, it can have a more fundamental role in quantum gravity and cosmology. The resolution of questions regarding quantum gravity, in my viewpoint, lies in a fresh and better understanding of gauge theories (note that gravity is a gauge theory with diffeomorphisms as gauge symmetry). The sandwich quantization proposal should be viewed as an attempt in this direction. The idea is that each super-selection sector in the physical Hilbert space should be associated with different observers. In other words, the role of observer is carried or manifested in the Hilbert space she uses; see the discussions below on “gauge equivalence principle”.
We stressed the already known, but often overlooked, fact that (e.g. see [5, 13, 15]) for quantization it is sufficient to impose the EoM for the gauge d.o.f, which appear as constraints, through sandwich conditions (8). We briefly discussed how the sandwich conditions may be solved, leading to a new set of super-selection sectors in the physical Hilbert space besides the textbook standard Class 1 states. The Class 1 physical Hilbert space, obtained by imposing gauge-fixing constraints through the right-action (6), is a direct extension of the classical constraints to the quantum level. The sandwich conditions and Class 2 solutions are the options arising through the quantization procedure, with no direct classical counterpart. The latter may be witnessed by the appearance of $\hbar$ in the right-hand side of (13). At the quantum level, we have the option that, while $\bm{\mathcal{C}}^{i}$ vanish when inserted into the $n$-point function of any generic local operators, the operator $\bm{\mathcal{C}}^{i}$ need not be well-defined over the physical Hilbert space $\mathcal{H}_{\text{p}}$. In other words, requiring that $\bm{\mathcal{C}}^{i}$ are well defined over $\mathcal{H}_{\text{p}}$ restricts us to the Class 1 Hilbert space, whereas relaxing this seemingly unnecessary requirement, opens up the possibility of super-selection sectors and the Class 2 physical Hilbert spaces. There are some steps are remaining to complete the sandwich quantization proposal and show its consistency and sufficiency.
@@
In this work, for illustration purposes we focused on the “axial-type gauges”. Consistency of the proposal requires that it should work for any arbitrary gauge-fixing. For these cases one should work through introducing ghosts and check the BRST invariance, that the physical states satisfying the sandwich conditions are defined up to BRST equivalence classes (see e.g. [10] for discussion on BRST symmetry). That is, we need to show that our each super-selection sector in the Class 2 physical Hilbert spaces is BRST invariant; in other words, one need to work through a “sandwich version” of the BRST symmetry and analysis.
@@
We need to uncover the physical meaning of the new sectors in the physical Hilbert space, the super-selection sectors in the Class 2 part. First we note that, Class 2 states by definition, are orthogonal to the Class 1 states, and the gauge-invariant dynamics of the theory will never mix them with the Class 1 states. That is, one can consistently formulate the theory using only the Class 1 states and recover the standard QFT textbook results. There is, however, another option: physics may be formulated based on either of the super-selection sectors in the Class 2 Hilbert space. Based on the example explored in [2] and the Maxwell theory example discussed in the previous section, we propose the “sandwich equivalence principle”:
Physical observables of gauge field theories should equivalently be described by the Class 1 or either of super-selection sectors in the Class 2 physical Hilbert space.
The above is somewhat similar to what we have in Einstein’s Equivalence Principle and the choice of different observers: each super-selection sector in the Class 2 physical Hilbert space provides states above a vacuum state associated with a particular physical observer. Of course, it remains to fully establish and formulate this equivalence principle. Conversely and recalling that Einstein’s general relativity is the gauge theory with diffeomorphism as the gauge symmetry, the gauge equivalence principle stated above may be viewed as a proposal for taking Einstein’s equivalence principle to quantum level. Noting that fixing diffeomorphisms is physically corresponding to choosing an observer, the “gauge equivalence principle” highlights the role of observer in quantum gravity and provides a framework to formulate it.